Find the equation of the normal to the solid 2x
2 −y
2 +8z
2 = 11 at a point where it
intersects the line x−3 = z =
y+1
2
.
1
Expert's answer
2020-03-02T15:25:32-0500

Solid: "f(x,y,z)=2x^2 \u2212y^2 +8z^2 - 11=0"

Line: "x−3 = z = \frac{y+1}{2}"

Determine the point of intersection.


"\begin{matrix}
 2x^2 −y^2 +8z^2 & = &11 \\
 z & = &x-3 \\
 z &= &\frac{y+1}{2}
\end{matrix}"


"\begin{matrix}
 2x^2 −y^2 +8z^2 & = &11 \\
 z & = &x-3 \\
 y &= &2z-1
\end{matrix}"


"\begin{matrix}
 2x^2 −y^2 +8z^2 & = &11 \\
 z & = &x-3 \\
 y &= &2(x-3)-1
\end{matrix}"

"\begin{matrix}
 2x^2 −y^2 +8z^2 & = &11 \\
 z & = &x-3 \\
 y &= &2x-7
\end{matrix}"


Substitute "z=x-3 \text{ and } y=2x-7" into the first equation.


"2x^2 - (2x-7)^2 + 8(x-3)^2=11"


Clear the parentheses and simplify.


"3x^2 - 10x + 6 =0"


Using the formula for the discriminant, obtain the solution of this equation.


"x_1= \frac{5}{3} + \frac{\sqrt{7}}{3}, \quad x_2= \frac{5}{3} - \frac{\sqrt{7}}{3}"


Determine the "y \text{ and }z."

"z_{1} = \frac{5}{3} +\frac{\sqrt{7}}{3} - 3 = -\frac{4}{3} + \frac{\sqrt{7}}{3}, \quad z_{2} = -\frac{4}{3} - \frac{\sqrt{7}}{3} \\

y_{1} = 2( -\frac{4}{3} + \frac{\sqrt{7}}{3}) - 1 = -\frac{11}{3} + \frac{2\sqrt{7}}{3}, \quad y_{2} = -\frac{11}{3} -\frac{2\sqrt{7}}{3}"


Thus, obtain two points of intersection "P_{1} = \left( \frac{5}{3} + \frac{\sqrt{7}}{3}, -\frac{11}{3} + \frac{2\sqrt{7}}{3}, -\frac{4}{3} + \frac{\sqrt{7}}{3}\right) \text{ and } P_{2} = \left( \frac{5}{3} - \frac{\sqrt{7}}{3}, -\frac{11}{3} - \frac{2\sqrt{7}}{3}, -\frac{4}{3} - \frac{\sqrt{7}}{3}\right)."


The equation of normal to solid at the point "P_{1}" is


"\frac{x-x_1}{f_{x}(x_1,y_1,z_1)} = \frac{y-y_1}{f_{y}(x_1,y_1,z_1)} = \frac{z-z_1}{f_{z}(x_1,y_1,z_1)}"

where "f_{x}, f_y, f_z" are the partial derivatives of "f" with respect to "x,y,z" accordingly. Determine the partial derivatives.

"f_x = 4x \\
f_y = -2y\\
f_z = 16z"


Therefore, the equation of the normal to the solid at the first point is



"\frac{x-\frac{5}{3}-\frac{\sqrt{7}}{3}}{4\cdot(\frac{5}{3}+\frac{\sqrt{7}}{3})}= \frac{y+\frac{11}{3}-\frac{2\sqrt{7}}{3}}{-2\cdot(-\frac{11}{3}+\frac{2\sqrt{7}}{3})} = \frac{z+\frac{4}{3}-\frac{\sqrt{7}}{3}}{16\cdot(-\frac{4}{3}+\frac{\sqrt{7}}{3})}"

and the equation of the normal to the solid at the second point is



"\frac{x-\frac{5}{3}+\frac{\sqrt{7}}{3}}{4\cdot(\frac{5}{3}-\frac{\sqrt{7}}{3})}= \frac{y+\frac{11}{3}+\frac{2\sqrt{7}}{3}}{-2\cdot(-\frac{11}{3}-\frac{2\sqrt{7}}{3})} = \frac{z+\frac{4}{3}+\frac{\sqrt{7}}{3}}{16\cdot(-\frac{4}{3}-\frac{\sqrt{7}}{3})}"